If the population in a region is around 117 million people now and if the population grows continuously at a relative growth rate of 1.45%, what will the population be in 7 years? Compute the answer to three significant digits.
1
Expert's answer
2012-06-14T08:35:17-0400
we use exponential model n(t)=n(0)*e^(0.0145*t)
so
n(7)=117*e^(0.0145*7)=129.499 million people